11^3 - 10^3 - 10^3 + 9^3 = ?
Is there an easier way to solve this?

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

11^3 - 10^3 - 10^3 + 9^3 = ?
Is there an easier way to solve this?
==================
No variable, nothing to solve.
11^3 - 2*10^3 + 9^3
= 1331 - 2000 + 729
= 60

Answer by rothauserc(4718)

11^3 - 10^3 - 10^3 + 9^3 =
:
11^3 -10^2(10+10) + (3^2)^3 =
:
11^3 - 10^2(20) + 3^6 =
:
11^3 - 2000 + 729 =
:
1331 - 2000 + 729 = 60
: